OpenStudy (anonymous):

(1.0, 228,000), (3.0, 204,000) If I have these numbers to work with, how do I figure out the slope-intercept

OpenStudy (amistre64):

( x , y ) (3.0, 204000) -(1.0, 228000) ------------- ( 2 , -24000 ) slope = y/x = -12000

OpenStudy (anonymous):

That is what I got, but I am having trouble with the slope-intercept of the problem it does want to work

OpenStudy (anonymous):

Can you tell me what numbers to plug in where. y=m(x)+b

OpenStudy (amistre64):

if my computer behaves :)

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

y = mx + b | | | y = -12000 x + 240000

OpenStudy (anonymous):

where did the 240,000 come from

OpenStudy (amistre64):

i plugged in an easy pointt from above: (1,228000) 228000 = -12000(1) + B 228000 +12000 -------- 240000 = B

OpenStudy (anonymous):

thank you